# TFusionOcc: Student’s t-Distribution Based Object-Centric Multi-Sensor Fusion Framework for 3D Occupancy Prediction

### v1

@@ Abstract
3D semantic occupancy prediction enables autonomous vehicles (AVs) to perceive fine-grained geometric and semantic structure of their surroundings from onboard sensors, which is essential for safe decision-making and navigation. Recent models for 3D semantic occupancy prediction have successfully addressed the challenge of describing real-world objects with varied shapes and classes. However, the intermediate representations used by existing methods for 3D semantic occupancy prediction rely heavily on 3D voxel volumes or a set of 3D Gaussians, hindering the model’s ability to efficiently and effectively capture fine-grained geometric details in the 3D driving environment. This paper introduces TFusionOcc, a novel object-centric multi-sensor fusion framework for predicting 3D semantic occupancy. By leveraging multi-stage multi-sensor fusion, Student’s t-distribution, and the T-Mixture model (TMM), together with more geometrically flexible primitives, such as the deformable superquadric (superquadric with inverse warp), the proposed method achieved state-of-the-art (SOTA) performance on the nuScenes benchmark. In addition, extensive experiments were conducted on the nuScenes-C dataset to demonstrate the robustness of the proposed method in different camera and lidar corruption scenarios. The code will be available at: https://github.com/DanielMing123/TFusionOcc
@@ I Introduction
Accurate and efficient modeling of the scene around autonomous vehicles (AV) is essential for safe and collision-free navigation. As technology advances, the introduction of a 3D semantic occupancy representation has successfully addressed the limitations of traditional 3D object detection and 2D bird’s-eye-view (BEV) segmentation networks, particularly for detecting irregular objects and predicting occupied space in 3D. Thus, 3D semantic occupancy provides a suitable representation of the vehicle’s surrounding environment. This representation inherently ensures geometric consistency of the AV’s driving scene and can accurately describe occluded areas. In addition, a multi-sensor fusion technique for 3D semantic occupancy prediction further enhances an AVs’ ability to model the 3D world.
Recent advances in voxel-based multi-sensor fusion methods [18, 21, 35, 16, 33, 4, 26] have achieved remarkable progress by integrating complementary information from different sensors, such as surround-view cameras, surround-view radars, and lidars. Demonstrating superior robustness and overall performance in several challenging light and weather conditions, these methods heavily rely on volumetric grid calculations (Fig 1, voxel-based approach), which hinders real-time application and makes them unsuitable for edge device deployment. Meanwhile, recent advances in object-centric multi-sensor fusion methods [38, 23] aim to address the limitations of the voxel-based approach by utilizing a set of 3D Gaussian primitives to only model the occupied area in the driving scene (Fig 1, 3D Gaussian-based approach), reducing the computational overhead caused by empty regions by concentrating representational capacity where it matters most. However, their modeling performance is inferior to that of voxel-based approaches. This performance gap is mainly due to the limited geometric shape representation capability of primitives and their weaker robustness to outliers.
@@
Figure 1: Pipeline of three approaches: Voxel-based approach (top-left), 3D-Gaussian-Primitive-based object-centric approach (top-right), and our approach (bottom-left).
To address the aforementioned limitations, we propose TFusionOcc (Fig 1, our approach), a novel Student’s t-distribution-based object-centric multi-sensor fusion framework. We adopt a multi-stage feature-level fusion strategy, including early-stage, middle-stage, and late-stage fusion, to better integrate complementary information across modalities and mitigate the drawbacks of each modality. For object-centric modeling, we adopt a Student’s t-distribution and the T-Mixture model (TMM) as a probability kernel, combined with several types of primitives, such as the general T-Primitive, superquadric, and deformable superquadric (superquadric with inverse-warp), to achieve greater geometric shape representation flexibility and robustness against outliers. The extensive experiments conducted on nuScenes [2] and nuScenes-C [31] datasets demonstrate that our proposed approach not only achieves state-of-the-art (SOTA) performance but also exhibits superior robustness against several sensor corruptions. The main contributions of this paper are summarized as follows:
  • We propose TFusionOcc, a novel object-centric multi-stage multi-sensor fusion framework that leverages different types of T primitives to efficiently and effectively model the 3D driving scene for 3D semantic occupancy prediction.
  • We propose a multi-gate cross-attention fusion module (MGCAFusion), which consists of a weighted summation and gated concatenation fusion modules, to combine the multi-modal features in an effective way.
  • We propose a set of Student’s t-distribution-based primitives, including the general T-Primitive, the T-Superquadric primitive, and the T-Superquadric with inverse warp primitive, to effectively model 3D driving scenes.
  • We thoroughly investigated the robustness of the proposed method with respect to the impact of different perception ranges and different lidar and camera corruption scenarios.
@@ II-A Voxel-Based Approach
In OccFusion [18], the authors extract a set of 3D voxel volumes from each different modality sensor, such as surround-view cameras, surround-view radars and lidar, followed by the feature fusion operation through the proposed dynamic feature fusion module to perform multi-sensor fusion-based 3D semantic occupancy prediction. The model exhibits strong robustness under challenging rain and night scenarios. In Co-Occ [21], the authors leverage the volume rendering technique, which originally derived from NeRF [15], to refine the intermediate 3D feature volumes. They also proposed a GSFusion module for efficient multimodal feature fusion, achieving SOTA performance. In OccCylindrical [16], the authors proposed a multi-sensor fusion-based framework that uses a cylindrical partition to perform voxelisation of a lidar-captured 3D point cloud and a pseudo 3D point cloud predicted by surround-view cameras, resulting in better preservation of 3D geometric information. The feature fusion and further refinement operations are performed under the cylindrical coordinate system. In FusionOcc [35], the authors proposed a cross-modality fusion module to generate a high-quality depth-aware visual feature, followed by a 2D to 3D view transformation, resulting in a depth-aware visual feature volumes under the Cartesian coordinate system. The lidar branch conducts voxelization and voxel encoding, resulting in a lidar feature volume also under the Cartesian coordinate system. Different modality feature volumes were fused via feature channel concatenation and further refined by a 3D Encoder. In DAOcc [33], the authors leverage multi-task learning to introduce an additional 3D supervision signal to guide intermediate features, leading to robust generalization and overall SOTA performance. Similarly, in Inverse++ [19], the authors leverage multi-task learning by incorporating an extra auxiliary 3D object detection branch. In addition, SDGOcc [4] incorporates a multi-task head to introduce a 2D semantic mask into the model training procedure, leveraging an additional 2D semantic mask to enhance the precision and densification of depth maps obtained from lidar point cloud projection, thereby boosting the model’s overall performance.
@@
II-B Object-Centric Modelling Approach
In the GaussianFormer series [9, 7], the authors proposed a 3D Gaussian-based object-centric modeling approach to improve computational efficiency. They further proposed a Gaussian-mixture model (GMM) to improve the model’s robustness. The 3D-to-3D Gaussian splatting operator, initially proposed in [10] for 3D-to-2D splatting, is further proposed to boost the model’s efficiency. In QuadricFormer [40], the authors proposed leveraging a superquadric primitive combined with a Gaussian probability kernel to achieve object-centric modeling and benefit from the geometric shape representation’s flexibility, thereby improving model performance. In GaussianFormer3D [38], the authors extend the GaussianFormer work into a multi-sensor fusion scenario. By utilizing a set of 3D Gaussian primitives as an intermediate feature representation, the authors designed a pipeline that enables each 3D Gaussian primitive to aggregate features from different modalities and fuse them to perform 3D primitive-to-3D voxel volume Gaussian splatting. Benefiting from the object-centric modeling technique and multi-modal fusion, the model achieves significant performance and efficiency improvements. In GaussianFusionOcc [23], the authors proposed a GaussianFusionBlock module that performs efficient and high-quality feature aggregation and fusion across different modalities, allowing high-quality 3D Gaussian primitive modeling and leading to SOTA performance.
Though our proposed approach belongs to the object-centric modeling category, the key difference of our work compared to previous ones is that we explored Student’s t-distribution and T-mixture model to enhance the model robustness. Furthermore, we proposed a multi-stage feature fusion pipeline to mitigate the drawbacks of each modality, resulting in better fused features. Lastly, we enhanced the primitive’s geometric flexibility to provide a better intermediate representation, thereby achieving SOTA performance.
@@ III TFusionOcc
where $F$ refers to a multi-sensor fusion framework, $P^{T}$ refers to a set of primitives based on T-distribution $P^{T}=\left\{p^{T}_{1},p^{T}_{2},...,p^{T}_{K}\right\}$, along with a set of query vectors $Q=\left\{q_{1},q_{2},...,q_{K}\right\}$, $P^{T}_{refined}$ refers to a set of T-distribution-based primitives refined through multi-sensor fusion framework, $N$ refers to the total number of surround-view images and $Splat\left(\cdot\right)$ refers to the splatting operation. In the final predicted 3D semantic occupancy $Occ\in R^{\left\{X\times Y\times Z\right\}}$, each grid is assigned a semantic property ranging from 0 to $C$, where $C$ refers to the total amount of semantic classes. In our case, a class value of 0 corresponds to an empty grid.
@@ III-A Overview
Figure 2: Overall architecture of TFusionOcc. The pipeline comprises two different modality branches and a multi-stage feature fusion branch. The camera branch extracts multi-scale visual features and predicts a pseudo 3D point cloud from surround-view images. The pseudo 3D point cloud is further projected and cylindrically partitioned, resulting in camera-based, multi-scale, dense depth maps and a voxel volume defined under cylindrical coordinates. The lidar branch applies a cylindrical partition followed by a 3D encoder to extract the lidar feature. Meanwhile, the lidar point cloud is also projected to generate lidar-based, multi-scale, sparse depth maps. The feature fusion branch adopts a multi-stage fusion strategy to merge all outputs from the two-modality branch and leverages a proposed transformer to refine the T-primitives property through fused features.
@@
The overall architecture is shown in Figure 2. For the camera branch, given $\left\{img^{1},img^{2},...,img^{N}\right\}$ surround-view images, a 2D encoder and a decoder are leveraged to extract $V^{Cam}=\left\{\left\{V_{n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$ multi-scale visual features. Meanwhile, a DepthNet is leveraged, resulting in a camera-based pseudo 3D point cloud $Pt^{cam}$ and a multi-scale camera-based dense depth maps $V_{depth}^{cam}=\left\{\left\{V_{depth\_n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{D\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$, where $D$ refers to the total amount of depth-bins. The pseudo 3D point cloud $Pt^{cam}$ is further divided into cylindrical voxels, resulting in a camera-based 3D voxel volume $Vol^{Cam}_{Cylin}$ defined in cylindrical coordinates. For the lidar branch, given a 3D point cloud $Pt^{lidar}$ from lidar, the cylindrical partition is first applied, resulting in a lidar-based 3D voxel volume $Vol^{Lidar}_{Cylin}$. Then a 3D encoder (e.g., Cylinder3D [39]) is used to extract lidar features $V^{Lidar}$ in cylindrical coordinates. Meanwhile, $Pt^{Lidar}$ is projected onto each surround-view camera frame, resulting in sparse multi-scale lidar-based depth maps $V_{depth}^{lidar}=\left\{\left\{V_{depth\_n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{D\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$. Then, a multi-stage feature fusion strategy is adopted. For the camera branch side, $V_{depth}^{cam}\oplus V_{depth}^{lidar}$ results in multi-scale fused dense depth maps $V_{depth}^{fuse}$. Then $V_{depth}^{fuse}$ is encoded by a multi-layer perceptron (MLP) and added to $V^{Cam}$, resulting in multi-scale depth-aware visual features $F_{depth}^{Cam}$. In addition, an outer product operation $F_{depth}^{Cam}\otimes V_{depth}^{fuse}$ is conducted, resulting in a multi-scale lifted depth-aware visual feature $Vol^{Cam}_{depth}$. For the lidar branch side, each occupied voxel center of $Vol^{Lidar}_{Cylin}$ serves as an anchor to project onto $V^{Cam}$, thereby aggregating semantic information. The aggregated semantic information is further fused with the lidar feature $Vol^{Lidar}_{Cylin}$ through an early-fusion module, resulting in a semantic-aware lidar feature $Vol^{Lidar}_{sem}$. The projection operation begins with a transition from a cylindrical coordinate system to a Cartesian coordinate system, followed by a transition to a new coordinate frame. The $Vol^{Cam}_{Cylin}$ and $Vol^{Lidar}_{Cylin}$ are fused through the skeleton merge module, resulting in a fused voxel volume $Vol^{Fused}_{Cylin}$ under cylindrical coordinates. The $Vol^{Fused}_{Cylin}$ is then used as primitive anchors to initialize T-Primitives $P^{T}$ under cylindrical coordinates. The $P^{T}$, $Vol^{Cam}_{depth}$ and $Vol^{Lidar}_{sem}$ are fed to the transformer module, which stacks the $N\times$ blocks to obtain a set of refined T-Primitives. Lastly, the refined set of T-Primitives performs a splatting operation, resulting in the final 3D semantic occupancy grid.
III-B Image 2D Encoder and Decoder for Surround-View Images
The purpose of the image encoder is to capture both the spatial and semantic features of the surround-view images. These features serve as the foundation for the subsequent feature fusion procedure, and lastly, the task of 3D semantic occupancy prediction. In our approach, we first use a 2D backbone network (e.g. ResNet50[5]) to extract visual features at multiple scales $V^{Cam}_{Encode}=\left\{\left\{V_{n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$. Subsequently, these features are fused using feature-pyramid networks (FPN), resulting in visual features $V^{Cam}=\left\{\left\{V_{n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$ that have the same feature channel dimmension and resolutions that are $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$ of the input image resolution, respectively. The deeper visual feature, with a smaller resolution, contains more semantic information. In contrast, the relatively shallower visual feature, with larger resolutions, provides richer spatial details.
@@ III-B Image 2D Encoder and Decoder for Surround-View Images
III-C Camera-based Pseudo 3D Point Cloud and Multi-Scale Cam Dense Depth Maps Generation

### v2

# TFusionOcc: T-Primitive Based Object-Centric Multi-Sensor Fusion Framework for 3D Occupancy Prediction

@@ Abstract
The prediction of 3D semantic occupancy enables autonomous vehicles (AVs) to perceive the fine-grained geometric and semantic scene structure for safe navigation and decision-making. Existing methods mainly rely on either voxel-based representations, which incur redundant computation over empty regions, or on object-centric Gaussian primitives, which are limited in modeling complex, non-convex, and asymmetric structures. In this paper, we present TFusionOcc, a T-primitive-based object-centric multi-sensor fusion framework for 3D semantic occupancy prediction. Specifically, we introduce a family of Student’s t-distribution-based T-primitives, including the plain T-primitive, T-Superquadric and deformable T-Superquadric with inverse warping, where the deformable T-Superquadric serves as the key geometry-enhancing primitive. We further develop a unified probabilistic formulation based on the Student’s t-distribution and the T-mixture model (TMM) to jointly model occupancy and semantics, and design a tightly coupled multi-stage fusion architecture to effectively integrate camera and LiDAR cues. Extensive experiments on nuScenes show state-of-the-art performance, while additional evaluations on nuScenes-C demonstrate strong robustness under most corruption scenarios. The code will be available at: https://github.com/DanielMing123/TFusionOcc
@@ I Introduction
Accurate and efficient scene representation is essential for autonomous driving, since downstream planning and decision-making rely on a reliable understanding of both geometry and semantics. Compared with conventional 3D object detection and BEV segmentation, 3D semantic occupancy prediction provides a more complete and denser description of the surrounding environment, especially for irregular structures, free space, and partially occluded regions. Multi-sensor fusion further improves this representation by combining the complementary strengths of cameras and LiDAR.
@@
Figure 1: Pipeline of three approaches: Voxel-based approach (top), 3D-Gaussian-Primitive-based object-centric approach (middle), and our approach (bottom).
Recent progress in multi-sensor 3D semantic occupancy prediction has been largely driven by voxel-based fusion methods [18, 21, 37, 16, 35, 4, 26], which integrate complementary signals from cameras and LiDAR. Although these methods have achieved strong performance, they generally rely on dense volumetric processing (Fig. 1, voxel-based approach), leading to substantial redundant computation over empty regions that dominate outdoor driving scenes. This inefficiency has motivated the development of object-centric representations, where a compact set of learnable primitives is used to model occupied regions more selectively. Among object-centric approaches[39, 23], Gaussian-based primitives have recently shown promise for improving efficiency and compact representation (Fig 1, 3D Gaussian-based approach). However, their geometric modeling capability remains limited when faced with non-convex, asymmetric, and structurally irregular scene elements that are common in real driving environments. In addition, Gaussian-like formulations are less effective in handling outliers and long-tailed spatial distributions, which may reduce robustness under noisy or corrupted observations. As a result, existing object-centric occupancy frameworks still face a performance gap when compared with strong voxel-based methods, especially when an accurate geometric representation is required.
To address these limitations, we propose TFusionOcc (Fig 1, our approach), a T-primitive-based object-centric multi-sensor fusion framework for 3D semantic occupancy prediction. Specifically, we introduce a family of Student’s t-distribution-based T-primitives, including the plain T-primitive, T-Superquadric and a deformable T-Superquadric with inverse warping, where the latter serves as the key geometry-enhancing primitive. We further develop a unified probabilistic formulation based on Student’s t-distribution and the T-mixture model (TMM) to jointly model occupancy and semantics, and design a compact multi-stage fusion pipeline to effectively integrate camera and LiDAR cues. In this way, TFusionOcc improves geometric representation while preserving the efficiency advantage of object-centric modeling. Extensive experiments on nuScenes [2] show that TFusionOcc achieves state-of-the-art (SOTA) overall performance and consistently benefits from stronger primitive geometry. In particular, our results indicate that the flexibility of primitive shapes is more decisive than simply increasing the primitive count, highlighting the importance of geometry-aware primitive design. Additional evaluations on nuScenes-C [32] in different corruption settings further demonstrate strong robustness in many challenging cases. The main contributions of this paper are summarized as follows:
  • We propose a family of Student’s t-distribution-based T-primitives for object-centric 3D occupancy prediction, including the plain T-primitive, T-Superquadric, and a deformable T-Superquadric with inverse warping. In particular, the deformable T-Superquadric significantly improves geometric modeling flexibility, enabling better representation of non-convex and asymmetric structures.
  • We develop a unified probabilistic modeling framework based on Student’s t-distribution and the T-mixture model, through which both occupancy and semantics are jointly modeled. This design improves robustness to outliers and complex real-world noise while alleviating the limitations of Gaussian-based primitives in modeling long-tailed spatial distributions.
  • We design a tightly coupled multi-sensor fusion pipeline to support effective T-primitive learning. Specifically, it includes fused-depth-guided visual feature lifting, transformer-based T-primitive-centric refinement, and the GCWSFusion module for adaptive cross-modal interaction, allowing T-primitives to better exploit cross-modal cues for 3D semantic occupancy prediction.
@@ II-A Voxel-Based Approach
In OccFusion [18], the authors extract a set of 3D voxel volumes from each different modality sensor, such as surround-view cameras, surround-view radars and LiDAR, followed by the feature fusion operation through the proposed dynamic feature fusion module to perform multi-sensor fusion-based 3D semantic occupancy prediction. The model exhibits strong robustness under challenging rain and night scenarios. In Co-Occ [21], the authors leverage the volume rendering technique, which was originally derived from NeRF [15], to refine the intermediate 3D feature volumes. They also proposed a GSFusion module for efficient multimodal feature fusion, achieving SOTA performance. In FusionOcc [37], the authors proposed a cross-modality fusion module to generate a high-quality depth-aware visual feature, followed by a 2D to 3D view transformation, resulting in volumes of depth-aware visual features under the Cartesian coordinate system. The LiDAR branch conducts voxelization and voxel encoding, resulting in a LiDAR feature volume also under the Cartesian coordinate system. Different modality feature volumes were fused via feature channel concatenation and further refined by a 3D Encoder. In OccCylindrical [16], the authors proposed a multi-sensor fusion-based framework that uses a cylindrical partition to perform voxelisation of a LiDAR-captured 3D point cloud and a pseudo 3D point cloud predicted by surround-view cameras, resulting in better preservation of 3D geometric information. The feature fusion and further refinement operations are performed under the cylindrical coordinate system. In AdaptiveOcc [34], the author explored using an efficient octree data structure to preserve the fine-grained geometry of the 3D scene, achieving a good balance between computational complexity and prediction performance for the voxel-grid-based approach. In DAOcc [35], the authors leverage multi-task learning to introduce an additional 3D supervision signal to guide intermediate features, leading to robust generalization and overall SOTA performance. Similarly, in Doracamom [40], the authors developed a joint 3D detection and occupancy prediction pipeline under a camera-radar fusion paradigm. By leveraging the strengths of multi-task learning and further incorporating temporal fusion, the proposed method achieved superior performance in both 3D object detection and 3D semantic occupancy prediction. In addition, SDGOcc [4] incorporates a multi-task head to introduce a 2D semantic mask into the model training procedure, leveraging an additional 2D semantic mask to enhance the precision and densification of depth maps obtained from LiDAR point cloud projection, thereby boosting the model’s overall performance.
@@
II-B Object-Centric Modeling Approach
In the GaussianFormer series [9, 7], the authors proposed an object-centric 3D Gaussian model together with a Gaussian mixture model (GMM) to improve computational efficiency and robustness. In GraphGSOcc [27], the authors combine the object-centric 3D Gaussian representation with a graph transformer to realize dynamic and static object decoupling modeling. Allow the model to optimize foreground dynamic objects and background static objects separately. The approach further incorporates a temporal fusion mechanism to achieve superior performance. In GaussianFormer3D [39], the authors extend the GaussianFormer work to a multi-sensor fusion scenario. Using a set of 3D Gaussian primitives as an intermediate feature representation, the authors designed a pipeline that enables each 3D Gaussian primitive to aggregate features from different modalities and fuse them to perform 3D primitive to-3D voxel volume Gaussian splatting. Benefiting from the object-centric modeling technique and multi-modal fusion, the model achieves significant performance and efficiency improvements. In GaussianFusionOcc [23], the authors proposed a GaussianFusionBlock module that performs efficient and high-quality feature aggregation and fusion across different modalities, allowing high-quality 3D Gaussian primitive modeling and leading to outstanding performance.
Unlike previous object-centric methods, our work introduces a family of Student’s t-distribution-based T-primitives for 3D semantic occupancy prediction. In particular, we propose a deformable T-Superquadric with inverse warping, which substantially strengthens primitive-level geometric flexibility. We further formulate occupancy and semantics under a unified TMM framework and design a multi-stage camera–LiDAR fusion pipeline to support effective T-primitive refinement. In this way, our method aims to preserve the efficiency advantage of object-centric modeling while improving geometric representation capability and robustness.
@@ III TFusionOcc
where $F$ refers to a multi-sensor fusion framework, $P^{T}$ refers to a set of primitives based on T-distribution $P^{T}=\left\{p^{T}_{1},p^{T}_{2},...,p^{T}_{K}\right\}$, along with a set of query vectors $Q=\left\{q_{1},q_{2},...,q_{K}\right\}$, $P^{T}_{refined}$ refers to a set of T-distribution-based primitives refined through multi-sensor fusion framework, $N$ refers to the total number of surround-view images and $Splat\left(\cdot\right)$ refers to the splatting operation. In the final predicted 3D semantic occupancy $Occ\in R^{\left\{X\times Y\times Z\right\}}$, each grid is assigned a semantic property ranging from 0 to $C$, where $C$ refers to the total number of semantic classes. In our case, a class value of 0 corresponds to an empty grid.
@@ III-A Overview
Figure 2: Overall architecture of TFusionOcc. The pipeline comprises a camera branch, a LiDAR branch, and a multi-stage fusion branch. The camera branch extracts multi-scale visual features and predicts a pseudo 3D point cloud from surround-view images. The pseudo 3D point cloud is further projected and cylindrically partitioned, resulting in camera-based, multi-scale, dense depth maps and a voxel volume defined under cylindrical coordinates. The LiDAR branch applies a cylindrical partition followed by a 3D encoder to extract the LiDAR feature. Meanwhile, the LiDAR point cloud is also projected to generate LiDAR-based, multi-scale, sparse depth maps. The feature fusion branch adopts a multi-stage fusion strategy to merge all outputs from the two-modality branch and leverages a proposed transformer to refine the T-primitives property through fused features.
@@
The overall architecture is shown in Fig. 2. For the camera branch, given $\left\{img^{1},img^{2},...,img^{N}\right\}$ surround-view images, a 2D encoder and a decoder are leveraged to extract $V^{Cam}=\left\{\left\{V_{n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{C\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$ multi-scale visual features. Meanwhile, a DepthNet is leveraged, resulting in a camera-based pseudo 3D point cloud $Pt^{cam}$ and a multi-scale camera-based dense depth maps $V_{depth}^{cam}=\left\{\left\{V_{depth\_n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{D\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$, where $D$ refers to the total amount of depth-bins. The pseudo 3D point cloud $Pt^{cam}$ is further divided into cylindrical voxels, resulting in a camera-based 3D voxel volume $Vol^{Cam}_{Cylin}$ defined in cylindrical coordinates. For the LiDAR branch, given a 3D point cloud $Pt^{lidar}$ from LiDAR, the cylindrical partition is first applied, resulting in a LiDAR-based 3D voxel volume $Vol^{Lidar}_{Cylin}$. Then a 3D encoder (e.g., Cylinder3D [41]) is used to extract LiDAR features $V^{Lidar}$ in cylindrical coordinates. Meanwhile, $Pt^{Lidar}$ is projected onto each surround-view camera frame, resulting in sparse multi-scale LiDAR-based depth maps $V_{depth}^{lidar}=\left\{\left\{V_{depth\_n}^{l}\right\}_{n=1}^{N}\in\mathbb{R}^{D\times H_{l}\times W_{l}}\right\}_{l=1}^{L=3}$. Then, a multi-stage feature fusion strategy is adopted. For the camera branch side, $V_{depth}^{cam}\oplus V_{depth}^{lidar}$ results in multi-scale fused dense depth maps $V_{depth}^{fuse}$. Then $V_{depth}^{fuse}$ is encoded by a multi-layer perceptron (MLP) and added to $V^{Cam}$, resulting in multi-scale depth-aware visual features $F_{depth}^{Cam}$. In addition, an outer product operation $F_{depth}^{Cam}\otimes V_{depth}^{fuse}$ is conducted, resulting in a multi-scale lifted depth-aware visual feature $Vol^{Cam}_{depth}$. For the LiDAR branch side, each occupied voxel center of $Vol^{Lidar}_{Cylin}$ serves as an anchor to project onto $V^{Cam}$, thereby aggregating semantic information. The aggregated semantic information is further fused with the LiDAR feature $Vol^{Lidar}_{Cylin}$ through an early-fusion module, resulting in a semantic-aware LiDAR feature $Vol^{Lidar}_{sem}$. The $Vol^{Cam}_{Cylin}$ and $Vol^{Lidar}_{Cylin}$ are fused through the skeleton merge module, resulting in a fused voxel volume $Vol^{Fused}_{Cylin}$ under cylindrical coordinates. The $Vol^{Fused}_{Cylin}$ is then used as primitive anchors to initialize T-Primitives $P^{T}$ under cylindrical coordinates. The $P^{T}$, $Vol^{Cam}_{depth}$ and $Vol^{Lidar}_{sem}$ are fed to the transformer module, which stacks the $N\times$ blocks to obtain a set of refined T-Primitives. Lastly, the refined set of T-Primitives performs a splatting operation, resulting in the final 3D semantic occupancy grid.
@@
III-B Camera-based Pseudo 3D Point Cloud and Multi-Scale Cam Dense Depth Maps Generation